# Quarkiton: a one-quark state near a boundary of confinement phase of QCD

### v1

@@ III.2 Extrapolation to zero temperature
We found that the free energy of the heavy quarks can be excellently described by the Cornell potential (11) in the whole range of studied temperatures, $0.507T_{c}<T<0.996T_{c}$. In particular, the long-distance tail of the Cornell potential (11) corresponds to a linear “stringy” potential that attracts the quark to the wall. Our numerical data shows that the corresponding “quarkiton” string tension $\sigma_{Q|}$, presented in Fig. 4, is a diminishing function of temperature $T$.
Figure 4: The quarkiton string tension $\sigma_{Q|}$, which determines the long-distance linear attraction (11) between a single heavy quark $Q$ and a neutral chromometallic mirror, as a function of temperature $T$. The solid line represents the best fit of the data by function (12).
@@
$\mathcal{R}_{Q|}=\lim_{T\to T_{c}^{-}}\frac{\sigma_{Q|}(T)}{\sigma_{Q\bar{Q}}(T)}=0.703(5)\,.$ (13)
However, this method could be criticized as inexact since the low-temperature plateau, shown in Fig. 4, is not sufficiently developed in the sense that there are not enough data points at the plateau. Unfortunately, below the lowest studied temperature, the data for the expectation value of the Polyakov loop becomes very noisy, thus prohibiting us from studying more points at the suspected plateau.
Figure 5: The ratio of the quarkitonic string tension to the fundamental string tension, $\sigma_{Q|}(T)/\sigma_{Q\bar{Q}}(T)$, for different temperatures. The average of the plateau in the range of temperatures $T\leqslant 0.75T_{c}$ is shown by the solid line, with the uncertainty shown by the shaded region.
@@
Figure 6: The Coulomb coefficient $\alpha_{Q|}$ in the Cornell potential (11) between a test quark and the mirror is shown as a function of the temperature $T$. The average of the plateau in the range of temperatures $T\leqslant 0.75T_{c}$ is represented by the solid line, with the uncertainty shown by the shaded region.
We also examined the Coulomb coefficient $\alpha_{Q|}$ in the Cornell potential (11) at various temperatures. The results, shown in Fig. 6, imply that the Coulomb coupling rises as a function of temperature at $T\gtrsim 0.75T_{c}$. At lower temperatures, the data are forming a plateau, as expected. The average of the Coulomb coefficient below the temperature $T=0.75T_{c}$ gives us the low-temperature limit $\alpha_{Q|}(T\to 0)=0.066(3)$.
@@ Acknowledgements.
The numerical simulations were performed on the equipment of Shared Resource Center "Far Eastern Computing Resource" IACP FEB RAS IAC and the computing cluster Vostok-1 of Far Eastern Federal University. The work of MNC was funded by the EU’s NextGenerationEU instrument through the National Recovery and Resilience Plan of Romania - Pillar III-C9-I8, managed by the Ministry of Research, Innovation and Digitization, within the project entitled “Facets of Rotating Quark-Gluon Plasma” (FORQ), contract no. 760079/23.05.2023 code CF 103/15.11.2022. The work of VAG and AVM was supported by Grant No. FZNS-2024-0002 of the Ministry of Science and Higher Education of Russia.

### v2

@@ III.2 Extrapolation to zero temperature
We found that the free energy of the heavy quarks can be excellently described by the Cornell potential (11) in the whole range of studied temperatures, $0.507T_{c}<T<0.996T_{c}$. In particular, the long-distance tail of the Cornell potential (11) corresponds to a linear “stringy” potential that attracts the quark to the wall. Our numerical data show that the corresponding quarkiton string tension $\sigma_{Q|}$, presented in Fig. 4, is a diminishing function of temperature $T$.
Figure 4: Quarkiton string tension $\sigma_{Q|}$, which determines the long-distance linear attraction (11) between a single heavy quark $Q$ and a neutral chromometallic mirror, as a function of temperature $T$. The solid line represents the best fit of the data by function (12).
@@
$\mathcal{R}_{Q|}=\lim_{T\to 0}\frac{\sigma_{Q|}(T)}{\sigma_{Q\bar{Q}}(T)}=0.703(5)\,.$ (13)
However, this method could be criticized as inexact since the low-temperature plateau, shown in Fig. 4, is not sufficiently developed in the sense that there are not enough data points at the plateau. Unfortunately, below the lowest studied temperature, the data for the expectation value of the Polyakov loop become very noisy, thus prohibiting us from studying more points at the suspected plateau.
Figure 5: Ratio of the quarkitonic string tension to the fundamental string tension, $\sigma_{Q|}(T)/\sigma_{Q\bar{Q}}(T)$, for different temperatures. The average of the plateau in the range of temperatures $T\leqslant 0.75T_{c}$ is shown by the solid line, with the uncertainty shown by the shaded region.
@@
Figure 6: Coulomb coefficient $\alpha_{Q|}$ in the Cornell potential (11) between a test quark and the mirror is shown as a function of the temperature $T$. The average of the plateau in the range of temperatures $T\leqslant 0.75T_{c}$ is represented by the solid line, with the uncertainty shown by the shaded region.
We also examined the Coulomb coefficient $\alpha_{Q|}$ in the Cornell potential (11) at various temperatures. The results, shown in Fig. 6, imply that the Coulomb coupling rises as a function of temperature at $T\gtrsim 0.75T_{c}$. At lower temperatures, the data form a plateau, as expected. The average of the Coulomb coefficient below the temperature $T=0.75T_{c}$ gives us the low-temperature limit $\alpha_{Q|}(T\to 0)=0.066(3)$.
@@ IV Conclusions
The numerical simulations were performed on the equipment of Shared Resource Center "Far Eastern Computing Resource" IACP FEB RAS IAC and the computing cluster Vostok-1 of Far Eastern Federal University.
@@ Acknowledgements.
The work of M.N.C. was funded by the EU’s NextGenerationEU instrument through the National Recovery and Resilience Plan of Romania-Pillar III-C9-I8, managed by the Ministry of Research, Innovation and Digitization, within the project entitled “Facets of Rotating Quark-Gluon Plasma” (FORQ), Contract No. 760079/23.05.2023 code CF 103/15.11.2022. The work of V.A.G. and A.V.M. was supported by Grant No. FZNS-2024-0002 of the Ministry of Science and Higher Education of Russia.
Data Availability Statement
The data that support the findings of this article are openly available Chernodub et al. (2025).